step by step explanation on solving the following equation please:
7 + 8x - 5 = 5x + 9 + 4x

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39616)

7 + 8x - 5 = 5x + 9 + 4x
-
Addition Property of Equality
Arithmetic & Factorization
Multiplication Property of Equality

x=-7

Answer by ikleyn(52776)

.

1.  Combine the like terms in each side. You will get

        8x + 2 = 9x + 9.



2.  Subtract 8x from both sides.  You will get

        2 = x + 9.



3.  Subtract 9 from both sides.  You will get

        -7 = x.



4.  So, x= -7 is the solution and is your answer.



ANSWER.  x= -7.

The solution is just completed.

--------------------

If you need to learn on how to solve equations (starting from the very beginning), look into the lesson
    - HOW TO solve a linear equation
in this site.